(5 points) Find area enclosed by f(x) = x+ + 128 and g(x) = 160 – x4. Answer:
Answer #1

fCm ) = x9 +128 a cm) 2160-24 Assea enclosed by f(M) and g(n) dy da x=-2 g =f(x) y=160-X 4 I dy da az-2 y 239 +178 160-24 da—100+ 一 50 150